# FoodLMM: A Versatile Food Assistant using Large Multi-modal Model

### v1

@@ Abstract
Large Multi-modal Models (LMMs) have made impressive progress in many vision-language tasks. Nevertheless, the performance of general LMMs in specific domains is still far from satisfactory. This paper proposes FoodLMM, a versatile food assistant based on LMMs with various capabilities, including food recognition, ingredient recognition, recipe generation, nutrition estimation, food segmentation and multi-round conversation. To facilitate FoodLMM to deal with tasks beyond pure text output, we introduce a series of novel task-specific tokens and heads, enabling the model to predict food nutritional values and multiple segmentation masks. We adopt a two-stage training strategy. In the first stage, we utilize multiple public food benchmarks for multi-task learning by leveraging instruct-following paradigm. In the second stage, we construct a multi-round conversation and a reasoning segmentation datasets to fine-tune the model, enabling it to conduct professional dialogues and generate segmentation masks based on complex reasoning in food domain. Our fine-tuned FoodLMM achieves state-of-the-art results across several food benchmarks. We will make our code, models and datasets publicly available.
[…]
@@ 1 Introduction
Although previous LMMs perform well on general images and questions, due to lack of domain-specific expertise, they often cannot provide reliable assistance in vertical domains and even produce incorrect responses or hallucinations [50]. This is particularly evident in the food realm. For instance, when asked about nutrition in a food image, general LMMs typically can only answer what nutritional elements it contains but fail to provide specific quantities and precise nutritional content of the food. In food domain, the previous efforts have been devoted to several tasks, ranging from food and ingredient recognition [3, 6, 2, 8, 97, 30], recipe generation [79, 80, 32, 86, 87], nutritional estimation [82] to food segmentation [28, 72, 95, 102, 22, 35, 49, 81]. Though the existing works achieve promising results in each individual task, they are incapable of dealing with other tasks using a single model.
@@
In this paper, we address the limitations by proposing FoodLMM, a versatile food assistant using Large Multi-modal model tailored for various food-related tasks, As shown in Figure 1. As far as we know, FoodLMM is the first unified and multitask LMM in the food domain. FoodLMM is capable of handling a variety of food-related tasks including: Food Classification, Ingredient Recognition, Recipe Generation, Nutrition Estimation, Referring Segmentation, and Reasoning Segmentation achieving state-of-the-art (SOTA) performance in each task. The architecture of FoodLMM extends from LISA [48], which consists of a base multi-modal large language model LLaVA [58], a segmentation model SAM [44] and a series of novel task-specific head tokens. Specifically, the outputs for three tasks including food classification, ingredient recognition and recipe generation are pure text and generated similarly to other LMMs [48, 50]. Importantly, we introduce multiple special segmentation tokens to the vocabulary and feed their hidden states to the segmentation decoder to generate one or multiple masks. Notably, by employing this mask generation paradigm with our elaborate instructions, FoodLMM is empowered with the capability to solve the most challenging one-to-many and one-to-zero problems [39] in the Referring Segmentation task. Additionally, a series of nutritional task-specific tokens are introduced for nutrition estimation, where their hidden states from the base LMM are processed through corresponding regression heads to yield precise values for various nutritional values.
@@
The training process of FoodLMM is divided into two stages. In the first stage, we utilize multiple public datasets [6, 79, 82, 95, 72] to conduct multi-task learning, aiming at injecting substantial basic food domain knowledge into FoodLMM. Specifically, we design a rich set of instructions and answer query templates for each task, enabling the model with the ability of handling different food tasks. The second stage aims to endow the model with multi-round conversational interaction capabilities in free-form prompts. For this purpose, we create a food multi-round conversation dataset FoodDialogues and a reasoning segmentation dataset FoodReasonSeg with GPT-4 [47], for training. The FoodDialogues is constructed based on the Nutriton5k food nutrition dataset, where detailed nutritional information is fed to GPT-4, prompting it to generate multi-round dialogues on various food-related topics, such as calorie calculation, dietary planning, and metabolism. Similarly, FoodReasonSeg is derived from the FoodSeg103 food segmentation dataset, utilizing the GPT-4 to create multi-round complex reasoning dialogues, the masks of the mentioned ingredients as reasoning segmentation labels. In the second training stage, these two datasets are added for fine-tuning. As a result, FoodLMM acquires the ability to interact in multi-round conversations with users and provide reasoning segmentation masks for complex queries. Our main contributions can be summarized as follows:
@@
  • We will release FoodDialogues and FoodReasonSeg datasets, along with our codebase and model checkpoints to facilitate future research.
@@ 2.1 Large Multi-modal Model
The remarkable natural language capabilities of large language models (LLMs) lead researchers to explore their use in visual tasks. Recently, GPT-4 [47] received significant recognition for introducing the ability to process multimodal inputs. Additionally, many multimodal LLMs that are available to the public have shown their effectiveness in general fields, such as Minigpt4 [101], Vicuna [18], LLaVA [58], MultiModal-GPT [31] and OpenFlamingo [1]. There are successful cases of applying large language models in areas such as biomedicine [90, 54, 60], finance [92], and law [19], but these applications are confined to a single text modality. LLaVA-Med [50] is a good specialization of LMM into the medical domain, which employs biomedical conversations generated by GPT-4 to instruct the model in learning biomedical knowledge. However, there are few applications of LMMs in other domains, especially the food domain, which motivate us to propose FoodLMM.
@@ 3 Food Visual Instruction-Following Data
$\mathbf{X}_{\text{system-message}}\color[rgb]{0.0,0.5,0.0}{\langle\text{STOP}\rangle}\color[rgb]{0,0,0}\backslash\text{n}$
@@
[…]
@@
FoodLMM’s training process consists of two stages. Stage 1 aims to to inject sufficient basic food knowledge into the LMM using various public food datasets. The conversational capability is empowered by our GPT-4 generated datasets in stage 2.
@@ 3.1 Stage 1: Public Food Datasets
Food VQA. We collectively refer to the three pure text output tasks: food classification, ingredient recognition and recipe generation as Food VQA and design instruction templates for each of them. We select two widely used benchmarks for Food VQA, VIREO Food-172 [6] and Recipe1M [79]. We only use recipes from the Recipe1M dataset that have corresponding images as cross-modal recipe retrieval [98, 6, 11].
Nutrition Estimation. Nutrition5k [82] provides fine-grained nutritional element values, food calories and quality annotations, including RGB images, multi-angle videos and depth image data. We select RGB images and video frames with good viewing angles as input images. Note that we do not use depth images, as they are difficult for users to acquire. We design seven types of instructions, including queries about overall calories or nutritional values, inquiries referring to one or more ingredients, questions about the primary nutrient, etc.
Food Segmentation. The FoodSeg103 [95] and UEC-FoodPIX Complete [72] datasets are the two most commonly used benchmarks for food segmentation, containing masks of 103 and 102 ingredients respectively. FoodSeg103 annotates dishes at a more fine-grained level and proves to be a more challenging benchmark for food image segmentation [49].
Figure 2: Architecture Overview. Food LMM consists of a base LMM (LLaVA [58]) with LoRA [36], a frozen segmentation encoder and a trainable segmentation decoder (SAM [44]), and a series of novel task-specific tokens for food segmentation and nutrition estimation.
@@ 3.2 Stage 2: GPT-4 Generated Conversation Datasets
FoodDialogues is built from the Nutrition5k dataset, which contains ingredient labels and precise nutrition information, making it unique and suitable for various conversational topics. Specifically, we follow the training and testing splits of the original data set and selected an overhead RGB image and a well-angled video frame for each sample. Send the sample’s ingredient list and detailed nutritional information to GPT-4 in the form of plain text, and request GPT-4 to generate multiple rounds of conversations on different topics, ranging from nutrition, calorie calculation, health and diseases, metabolism, dietary planning, allergies, food pairing to substitution. Answers are required to be professional and explanatory via text prompt for GPT-4.
@@ 4 Method
The network architecture of FoodLMM is illustrated in Figure 2. Our FoodLMM is built upon Large Multi-modal Model (LMM) LISA [48], which is based on LLaVA [50], taking image and text prompts as input. We adopted a two-stage training strategy to specialize the general LMM in the food domain. The first training stage aims to inject sufficient food-related basic knowledge into the LMM by using instruction-following data constructed in 3.1 based on five tasks in the food domain. As the output of the three tasks of food classification, ingredient recognition and recipe generation are pure text, these tasks could be learned through pure language auto-regressive similar to LLaVa. The output of food segmentation is multi-modal, with both masks and text, and nutrition estimation aims to produce nutrition values e.g., mass and calorie. We propose task-specific tokens and heads for food segmentation and nutrition estimation to address this issue. The segmentation masks are generated through multiple additional segmentation tokens and the segmentation encoder-decoder. Different types of nutritional values are predicted by a series of nutrition tokens and regression heads. The LMM is fine-tuned using LoRA to learn the generation of task-specific tokens and the answers for various tasks. The second training stage focuses on endowing FoodLMM with the capability to conduct multiple rounds of highly professional conversations and providing detailed explanations with segmentation masks for queries requiring complex reasoning.
@@ 4.1 Stage 1: Multi-task learning
The hidden states of these nutritional tokens are processed into embeddings through a Multilayer Perceptron (MLP). These embeddings are fed into the corresponding regression heads to get the predicted values. Finally, the predicted values replace the task-specific tokens in the original text output to obtain the answer with nutritional value. We employ the Mean Absolute Error (MAE) and the Mean Squared Error Loss (MSE) as follows:
@@
Referring Segmentation. To embed ingredient segmentation into our FoodLMM, we introduce segmentation tokens, denoted as $\langle seg_{i}\rangle$, with $i$ representing the ingredient’s index, into the primary vocabulary. When the model receives a query and its associated image, it generates a text response containing tokens for the ingredients to be segmented. The hidden states of these tokens are converted through a MLP into embeddings that capture the relevant ingredient information. These embeddings, combined with the visual representation from SAM’s encoder, are processed by the SAM decoder to create the ingredient masks.
@@
Almost all the referring segmentation models are limited to producing a single binary mask for different queries [37, 53, 61, 96] To address this issue, we aim to enable FoodLMM to go Beyond One-to-One [39], i.e., to produce different numbers of masks based on the queries. To achieve this, we design three different types of referring instructions: (1) to segment certain specified ingredients (one-to-one/one-to-many), e.g., Query: Segment the fish and lemon in this picture. Answer: The fish is masked as $\langle seg_{1}\rangle$ and the lemon as $\langle seg_{2}\rangle$. (2) to segment all visible ingredients (one-to-many), e.g., Query: Segment all the ingredients in this photo. Answer: The fish is masked as $\langle seg_{1}\rangle$, lemon as $\langle seg_{2}\rangle$, carrot as $\langle seg_{3}\rangle$ and green beans as $\langle seg_{4}\rangle$. (3) to segment objects that are not present in the image (one-to-zero), e.g., Query: Segment the watermelon in this picture. Answer: The watermelon is not found in this picture. Through these three instructions, the challenge of referring segmentation is dexterously addressed, realizing One-to-Any segmentation.
Our FoodLLM is trained by combining all the tasks and data with multitask learning. Thus, the overall loss $\mathcal{L}$ consists of $\mathcal{L}_{txt}$, $\mathcal{L}_{mask}$ and $\mathcal{L}_{nutrition}$, weighted by $\lambda_{txt}$, $\lambda_{nutrition}$ and $\lambda_{mask}$, as formulated below:
@@ 5.1 Experimental Setup
Training Details. We train our FoodLLM from LISA-7B-v1-explanatory model [48], using four NVIDIA 40G A100 GPUs. AdamW optimizer combined with the WarmupDecayLR is used as a learning rate scheduler. We set the initial learning rate to 0.0003, the weight decay to 0, and used 100 warm-up iterations. We set the weights $\lambda_{BCE}$ and $\lambda_{Dice}$ to 2.0 and 0.5 respectively, $\lambda_{MAE}$ and $\lambda_{MSE}$ to 0.1 and 0.0001 respectively, other weights are all set to 1. Throughout the training, we employed LoRA to fine-tune LLaVa. In addition, the parameters for SAM’s decoder, the MLP, and the heads dedicated to nutritional value estimation are all trainable.
In the stage 1, the model is trained with a batch size of 4 and the sampling rate of three tasks, i.e., Food VQA, Nutrition Estimation and Food Segmentation is set to $2:1:1$. We denote the model trained from Stage 1 as FoodLMM S1. In Stage 2, the batch size is reduced to $2$, and the sampling ratio of Multi-turn Conversation and Reasoning Segmentation is set to $3:2$. To retain the capabilities of the tasks in the first stage, we replayed the training data of the first stage, and the total sampling ratio of the generated data to the public data is set to $15:7$. The model trained from Stage 2 is denoted as FoodLMM Chat. To better evaluate the ability of FoodLMM on specific tasks, we fine-tuned FoodLMM on each task, recorded as FoodLMM FT, and the batch size during fine-tuning is set to 4.
@@
For food classification, top-1 accuracy is adopted as evaluation metric. While for ingredient recognition, the Intersection over Union (IoU) and F1-score are used to evaluate the performances. For recipe generation, we follow [80, 17] and apply SacreBLEU, Rouge-L metrics to quantify the quality of the generated recipes. For nutrition estimation, following [82], we use mean absolute error (MAE) and the percent of MAE to the respective mean for that field to measure regression accuracy for calories, mass, and individual macronutrient mass. Caloric MAE is measured in kilocalories, all others are measured in grams. For referring segmentation, in line with LISA [48], we employ the Complete Intersection over Union (cIoU) metric to evaluate the performance. Meanwhile, to evaluate the risk of our model incorrectly rejecting segmenting existing ingredients, we also report the probability of accurate response. For one-to-zero referring segmentation, we report the accuracy of our model successfully rejecting to segment non-existent items. For reasoning gegmentation, following LISA [48], Generalized Intersection over Union (gIoU) along with cIoU are adopted for the evaluation of Reasoning Segmentation.
@@
Table 3: Performance comparison on different food benchmarks. “FoodLMM S1” denotes the model trained from the first stage, and “FoodLMM FT” represents the model finetuned on specific tasks.

### v2

@@ Abstract
Large Multi-modal Models (LMMs) have made impressive progress in many vision-language tasks. Nevertheless, the performance of general LMMs in specific domains is still far from satisfactory. This paper proposes FoodLMM, a versatile food assistant based on LMMs with various capabilities, including food recognition, ingredient recognition, recipe generation, nutrition estimation, food segmentation and multi-round conversation. To facilitate FoodLMM to deal with tasks beyond pure text output, we introduce a series of novel task-specific tokens and heads, enabling the model to predict food nutritional values and multiple segmentation masks. We adopt a two-stage training strategy. In the first stage, we utilize multiple public food benchmarks for multi-task learning by leveraging the instruct-following paradigm. In the second stage, we construct a multi-round conversation dataset and a reasoning segmentation dataset to fine-tune the model, enabling it to conduct professional dialogues and generate segmentation masks based on complex reasoning in the food domain. Our fine-tuned FoodLMM achieves state-of-the-art results across several food benchmarks. We will make our code, models and datasets publicly available.
Keywords:
Food Assistant Large Multi-modal Models Multi-tasks
@@ 1 Introduction
Although previous LMMs perform well on general images and questions, due to a lack of domain-specific expertise, they often cannot provide reliable assistance in vertical domains and even produce incorrect responses or hallucinations [49]. This is particularly evident in the food realm. For instance, when asked about nutrition in a food image, general LMMs typically can only answer what nutritional elements it contains but fail to provide specific quantities and precise nutritional content of the food. In food domain, the previous efforts have been devoted to several tasks, ranging from food and ingredient recognition [3, 8, 2, 9, 99, 30], recipe generation [81, 80, 32, 87, 88], nutritional estimation [83] to food segmentation [28, 73, 96, 104, 23, 35, 48, 82]. Though the existing works achieve promising results in each individual task, they are incapable of dealing with other tasks using a single model.
@@
In this paper, we address the limitations by proposing FoodLMM, a versatile food assistant using a Large Multi-modal model tailored for various food-related tasks, As shown in Figure 1. To the best of our knowledge, FoodLMM is the first unified and multitask LMM in the food domain. FoodLMM is capable of handling a variety of food-related tasks including: Food Classification, Ingredient Recognition, Recipe Generation, Nutrition Estimation, Referring Segmentation, and Reasoning Segmentation, achieving state-of-the-art (SOTA) performance in each task. The architecture of FoodLMM extends from LISA [47], which consists of a base multi-modal large language model LLaVA [57], a segmentation model SAM [43] and a series of novel task-specific head tokens. Specifically, our FoodLMM produces text outputs for food classification, ingredient recognition and recipe generation tasks in an instruct-following fashion, similar to other LMMs [47, 49]. Importantly, we introduce multiple special segmentation tokens to the vocabulary and feed their hidden states to the segmentation decoder to generate one or multiple masks. Notably, by employing this mask generation paradigm with our elaborate instructions, FoodLMM is empowered with the capability to solve the most challenging one-to-many and one-to-zero problems [38] in the Referring Segmentation task. Additionally, a series of nutritional task-specific tokens are introduced for nutrition estimation, where their hidden states from the base LMM are processed through corresponding regression heads to yield precise values for various nutritional values.
@@
The training process of FoodLMM is divided into two stages. In the first stage, we utilize multiple public datasets [8, 81, 83, 96, 73] to conduct multi-task learning, aiming at injecting substantial basic food domain knowledge into FoodLMM. Specifically, we design a rich set of instructions and answer query templates for each task, enabling the model with the ability to handle different food tasks. The second stage aims to endow the model with multi-round conversational interaction capabilities in free-form prompts. For this purpose, we create a food multi-round conversation dataset FoodDialogues and a reasoning segmentation dataset FoodReasonSeg with GPT-4 [46], for training. The FoodDialogues is constructed based on the Nutriton5k food nutrition dataset, where detailed nutritional information is fed to GPT-4, prompting it to generate multi-round dialogues on various food-related topics, such as calorie calculation, dietary planning, and metabolism. Similarly, FoodReasonSeg is derived from the FoodSeg103 food segmentation dataset, utilizing the GPT-4 to create multi-round complex reasoning dialogues, the masks of the mentioned ingredients as reasoning segmentation labels. In the second training stage, these two datasets are added for fine-tuning. As a result, FoodLMM acquires the ability to interact in multi-round conversations with users and provide reasoning segmentation masks for complex queries. Our main contributions can be summarized as follows:
@@
  • We construct diverse multi-round conversation food dataset FoodDialogues and food reasoning segmentation dataset FoodReasonSeg. We will release both datasets, along with our codebase and model checkpoints to facilitate future research.
@@ 2.1 Large Multi-modal Model
The remarkable natural language capabilities of large language models (LLMs) lead researchers to explore their use in visual tasks. Recently, GPT-4 [46] received significant recognition for introducing the ability to process multimodal inputs. Additionally, many multimodal LLMs that are available to the public have shown their effectiveness in general fields, such as Minigpt4 [103], Vicuna [18], LLaVA [57], MultiModal-GPT [31] and OpenFlamingo [1]. There are successful cases of applying large language models in areas such as biomedicine [91, 53, 60], finance [93], and law [20], but these applications are confined to a single text modality. LLaVA-Med [49] is a good specialization of LMM into the medical domain, which employs biomedical conversations generated by GPT-4 to instruct the model in learning biomedical knowledge. However, there are few applications of LMMs in other domains, especially the food domain, which motivates us to propose FoodLMM.
@@ 3 Food Visual Instruction-Following Data
$\mathbf{X}_{\text{system-message}}\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}{\langle\text{STOP}\rangle}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\backslash\text{n}$
@@
USER: $\mathbf{X}^{2}_{\mathrm{q}}\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}{\langle\text{STOP}\rangle}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\backslash\text{n}\text{ASSISTANT: }\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}{\mathbf{X}^{2}_{\mathrm{a}}\langle\text{STOP}\rangle}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\backslash\text{n}\ldotsUSER$
@@
FoodLMM’s training process consists of two stages. Stage 1 aims to inject sufficient basic food knowledge into the LMM using various public food datasets. The conversational capability is empowered by our GPT-4 generated datasets in stage 2.
@@ 3.1 Stage 1: Public Food Datasets
Food VQA. We collectively refer to the three pure text output tasks: food classification, ingredient recognition and recipe generation as Food VQA and design instruction templates for each of them. We select two widely used benchmarks for Food VQA, VIREO Food-172 [8] and Recipe1M [81]. We only use recipes from the Recipe1M dataset that have corresponding images as cross-modal recipe retrieval [100, 8, 6]. We have elaborately designed 10, 8 and 11 instruction templates for food classification, ingredient recognition and recipe generation tasks respectively. Please refer to the supplementary materials for more details.
Nutrition Estimation. Nutrition5k [83] provides fine-grained nutritional element values, food calories and quality annotations, including RGB images, multi-angle videos and depth images. We select RGB images and video frames with good viewing angles. Note that we do not use depth images, as they are difficult for users to acquire. We design a rich collection of instruction templates for nutrient estimation tasks that contain 7 types of instructions, including queries about overall calories or nutritional values, inquiries referring to one or more ingredients, questions about the primary nutrient, etc. There are 64 query templates and 68 answer templates in total, listed in the supplementary material.
Food Segmentation. The FoodSeg103 [96] and UEC-FoodPIX Complete [73] datasets are the two most commonly used benchmarks for food segmentation, containing masks of 103 and 102 ingredients respectively. FoodSeg103 annotates dishes at a more fine-grained level and proves to be a more challenging benchmark for food image segmentation [48]. We design 10 query templates and 9 answer templates for food segmentation task. All the instruction templates can be found in the supplementary material.
Figure 2: Architecture Overview of the proposed FoodLMM.
@@ 3.2 Stage 2: GPT-4 Generated Conversation Datasets
FoodDialogues is built from the Nutrition5k dataset, which contains ingredient labels and precise nutrition information, making it unique and suitable for various conversational topics. Specifically, we follow the training and testing splits of the original data set and selected an overhead RGB image and a well-angled video frame for each sample. Send the sample’s ingredient list and detailed nutritional information to GPT-4 in the form of plain text, and request GPT-4 to generate multiple rounds of conversations on different topics, ranging from nutrition, calorie calculation, health and diseases, metabolism, dietary planning, allergies, food pairing to substitution. The prompt for GPT-4 is carefully designed based on expert opinions where answers are required to be professional and explanatory.
@@
Moreover, we conduct a human evaluation to measure the quality of the generated FoodDialogues and FoodReasonSeg datasets. Specifically, users are asked to rate the quality of GPT-4 generated content on a five-point scale (5 being the best). The average score is 4.2 for FoodDialogues and 4.4 for FoodReasonSeg, with more than half of the participants choosing 5, which showcases the remarkable reliability and high quality of the datasets. More details are provided in the supplementary material.
@@ 4 Method
The network architecture of FoodLMM is illustrated in Figure 2. Our FoodLMM takes image and text prompts as input, and is capable of handling various food tasks in a united model. The architecture of FoodLMM is built upon Large Multi-modal Model (LMM) LISA [47], which is LMM with powerful language generation and image segmentation ability. As we not only focus on language generation tasks, such as ingredient recognition and recipe generation but also food segmentation, we naturally choose LISA instead of other LMMs, such as LLaVA [49], to build our FoodLMM.
We adopted a two-stage training strategy to specialize the general LMM in the food domain. The first training stage aims to inject sufficient food-related basic knowledge into the LMM by using instruction-following data constructed in 3.1 based on five tasks in the food domain. As the output of the three tasks of food classification, ingredient recognition and recipe generation are pure text, these tasks could be learned through pure language auto-regressive similar to LLaVa. The output of food segmentation is multi-modal, with both masks and text, and nutrition estimation aim to produce nutrition values e.g., mass and calorie. We propose task-specific tokens and heads for food segmentation and nutrition estimation to address this issue. The segmentation masks are generated through multiple additional segmentation tokens and the segmentation encoder-decoder. Different types of nutritional values are predicted by a series of nutrition tokens and regression heads. The LMM is fine-tuned using LoRA to learn the generation of task-specific tokens and the answers for various tasks. The second training stage focuses on endowing FoodLMM with the capability to conduct multiple rounds of highly professional conversations and provide detailed explanations with segmentation masks for queries requiring complex reasoning.
@@ 4.1 Stage 1: Multi-task learning
The hidden states of these nutritional tokens are processed into embeddings through a Multilayer Perceptron (MLP). These embeddings are fed into the corresponding regression heads to get the predicted values. Finally, the predicted values replace the task-specific tokens in the original text output to obtain the answer with nutritional value.
We employ the Mean Absolute Error (MAE) and the Mean Squared Error Loss (MSE) as follows:
@@
Referring Segmentation. To embed ingredient segmentation into our FoodLMM, we introduce segmentation tokens, denoted as $\langle seg_{i}\rangle$, with $i$ representing the ingredient’s index, into the primary vocabulary. When the model receives a query and its associated image, it generates a text response containing tokens for the ingredients to be segmented. The hidden states of these tokens are converted through an MLP into embeddings that capture the relevant ingredient information. These embeddings, combined with the visual representation from SAM’s encoder, are processed by the SAM decoder to create the ingredient masks.
@@
Almost all the referring segmentation models are limited to producing a single binary mask for different queries [36, 52, 61, 97]. To address this issue, we aim to enable FoodLMM to go Beyond One-to-One [38], i.e., to produce different numbers of masks based on the queries. To achieve this, we design three different types of referring instructions: (1) to segment certain specified ingredients (one-to-one/one-to-many), e.g., Query: Segment the fish and lemon in this picture. Answer: The fish is masked as $\langle seg_{1}\rangle$ and the lemon as $\langle seg_{2}\rangle$. (2) to segment all visible ingredients (one-to-many), e.g., Query: Segment all the ingredients in this photo. Answer: The fish is masked as $\langle seg_{1}\rangle$, lemon as $\langle seg_{2}\rangle$, carrot as $\langle seg_{3}\rangle$ and green beans as $\langle seg_{4}\rangle$. (3) to segment objects that are not present in the image (one-to-zero), e.g., Query: Segment the watermelon in this picture. Answer: The watermelon is not found in this picture. Through these three instructions, the challenge of referring segmentation is dexterously addressed, realizing One-to-Any segmentation.
Our FoodLMM is trained by combining all the tasks and data with multitask learning. Thus, the overall loss $\mathcal{L}$ consists of $\mathcal{L}_{txt}$, $\mathcal{L}_{mask}$ and $\mathcal{L}_{nutrition}$, weighted by $\lambda_{txt}$, $\lambda_{nutrition}$ and $\lambda_{mask}$, as formulated below:
@@ 5.1 Experimental Setup
Training Details. We train our FoodLMM from LISA-7B-v1-explanatory model [47], using four NVIDIA 40G A100 GPUs. AdamW optimizer combined with the WarmupDecayLR is used as a learning rate scheduler. We set the initial learning rate to 0.0003, the weight decay to 0, and used 100 warm-up iterations. Following LISA, We set the weights $\lambda_{BCE}$ and $\lambda_{Dice}$ to 2.0 and 0.5 respectively. For the nutrition estimation task, we empirically adjust the weights $\lambda_{MAE}$ and $\lambda_{MSE}$ to 0.1 and 0.0001 respectively to balance their magnitudes with other losses. Throughout the training, we employed LoRA to fine-tune LLaVa. In addition, the parameters for SAM’s decoder, the MLP, and the heads dedicated to nutritional value estimation are all trainable.
In stage 1, the model is trained with a batch size of 4 and the sampling rate of three tasks, i.e., Food VQA, Nutrition Estimation and Food Segmentation is set to $2:1:1$. We denote the model trained from Stage 1 as FoodLMM S1. In Stage 2, the batch size is reduced to $2$, and the sampling ratio of Multi-turn Conversation and Reasoning Segmentation is set to $3:2$. To retain the capabilities of the tasks in the first stage, we replayed the training data of the first stage, and the total sampling ratio of the generated data to the public data is set to $15:7$. The model trained from Stage 2 is denoted as FoodLMM Chat. To better evaluate the ability of FoodLMM on specific tasks, we fine-tuned FoodLMM on each task, recorded as FoodLMM FT, and the batch size during fine-tuning is set to 4.
@@
For food classification, top-1 accuracy is adopted as the evaluation metric. While for ingredient recognition, the Intersection over Union (IoU) and F1-score are used to evaluate the performances. For recipe generation, we follow [80, 17] and apply SacreBLEU, Rouge-L metrics to quantify the quality of the generated recipes. For nutrition estimation, following [83], we use mean absolute error (MAE) and the percent of MAE to the respective mean for that field to measure regression accuracy for calories, mass, and individual macronutrient mass. Caloric MAE is measured in kilocalories, all others are measured in grams. For referring segmentation, in line with LISA [47], we employ the Complete Intersection over Union (cIoU) metric to evaluate the performance. Meanwhile, to evaluate the risk of our model incorrectly rejecting segmenting existing ingredients, we also report the probability of accurate response. For one-to-zero referring segmentation, we report the accuracy of our model successfully rejecting to segment non-existent items. For reasoning gegmentation, following LISA [47], Generalized Intersection over Union (gIoU) along with cIoU are adopted for the evaluation of Reasoning Segmentation.
@@
Table 3: Performance comparison on different food benchmarks. “FoodLMM S1" denotes the model trained from the first stage, and “FoodLMM FT" represents the model finetuned on specific tasks.